15x3(2)-2(3)=15 is my problem
i don't know how to arrive at this solution

Answer by checkley77(12844)

15x3(2)-2(3)=15
30x^3-6=15
x^3=15/30
x^3=1/2
x=cubert(1/2)
x=.7937 ans.